Calculate how many moles of NH3 form when each quantity of reactant completely reacts.
3N2H4(l)→4NH3(g)+N2(g)

2.7 mol N2H4
1
Expert's answer
2017-06-05T08:35:10-0400
Solution: n(NH3) = n(N2H4) *4/3 ; n(NH3) =(2.7*4)/3=3.6 (moles).
Answer: 3.6 moles.